### v1

# Partial identification of principal causal effects under violations of principal ignorability

Abstract
Principal stratification is a general framework for studying causal mechanisms involving post-treatment variables. When estimating principal causal effects, the principal ignorability assumption is commonly invoked, which we study in detail in this manuscript. Our first key contribution is studying a commonly used strategy of using parametric models to jointly model the outcome and principal strata without requiring the principal ignorability assumption. We show that even if the joint distribution of principal strata is known, this strategy necessarily leads to only partial identification of causal effects, even under very simple and correctly specified outcome models. While principal ignorability can lead to point identification in this setting, we discuss alternative, weaker assumptions and show how they lead to more informative partial identification regions. An additional contribution is that we provide theoretical support to strategies used in the literature for identifying association parameters that govern the joint distribution of principal strata. We prove that this is possible, but only if the principal ignorability assumption is violated. Additionally, due to partial identifiability of causal effects even when these association parameters are known, we show that these association parameters are only identifiable under strong parametric constraints. Lastly, we extend these results to more flexible semiparametric and nonparametric Bayesian models.
Principal stratification,
Principal ignorability,
Partial identification,
Bayesian inference,
keywords:
[…]
and
1 Introduction
In many causal inference problems, there exist post-treatment (also referred to as intermediate) variables, which can pose challenges for identification and inference. Principal stratification [11] is a general framework for studying causal mechanisms involving such post-treatment variables. It formally defines causal estimands of interest in terms of principal strata, where principal strata represent the joint potential values of the intermediate variable. The principal strata are not affected by the treatment and can therefore be treated as pre-treatment covariates, which leads to well-defined causal estimands conditional on the principal strata. These are local causal effects in the sense that they are defined for sub-populations of units defined by principal strata. Many scientific questions can be addressed within this framework, such as mediation [38, 30, 23], noncompliance [18, 31], surrogate endpoints [43], or truncation by death [41, 8, 40].
In settings with binary treatments, principal strata are represented by the joint values of potential intermediates under both treatment levels. At most we get to observe only one of these two potential intermediates, which means that we require additional structural assumptions in order to identify principal causal effects (PCEs). Monotonicity and exclusion restriction assumptions, which reduce the number of principal strata and place restrictions on causal effects in certain strata, are commonly used, particularly in settings with noncompliance in randomized trials [1, 16, 36, 3, 5]. Principal ignorability is another commonly invoked assumption [22, 19, 29], which enforces that the potential outcome under one treatment level only depends on the potential intermediate at that same treatment level. All of these assumptions are fundamentally not testable, and their plausibility should be judged within the context of the problem being studied.
Given the reliance on key untestable assumptions, many researchers have worked to develop approaches accounting for potential violations of causal assumptions. One approach is to perform sensitivity analysis, which highlights how results would change under specific violations of core assumptions, such as the principal ignorability assumption [9, 39, 33]. There is also an extensive literature on partial identification for principal causal effects [6, 17, 26, 31, 32, 40] where causal effects can be bounded under weaker assumptions than those required for point identification. Most of these previous studies have focused on randomized studies with binary intermediate variables. Others have focused on specific situations such as those with truncation by death, or when a secondary outcome is observed, which can provide more informative bounds. We focus on more general situations throughout this manuscript, or when such additional information is not available.
Another line of research aims to incorporate parametric assumptions to obtain identification when nonparametric identification is difficult or impossible to obtain. This is particularly the case with continuous intermediate variables, which we address in this manuscript, as standard assumptions such as monotonicity and exclusion restrictions are not sufficient for nonparametric identification. Recent works studying nonparametric identification in this more difficult scenario have made the principal ignorability assumption, along with the strong parametric assumption that the joint distribution of the two potential intermediates is governed by a known copula with a known correlation parameter [19, 28, 42]. Results are then estimated across a range of plausible correlation parameters to see if results are robust to this choice. Other works have attempted to estimate this unknown correlation parameter by jointly modeling the potential intermediates and observed outcomes simultaneously [37, 4]. The idea behind this strategy is that one can’t directly estimate the association between the two potential intermediates because they are never jointly observed, but this information can be obtained from a correctly specified outcome regression model. Related research aims to jointly model the distribution of the potential intermediates and observed outcome using either parametric [21] or semiparametric models [24, 25], and assume the correlation between the two potential intermediates is known. Many of these model-based approaches do not make the principal ignorability assumption, and they allow the observed outcome to depend on the values of both potential intermediate values.
In this article, we make a number of contributions to the literature on identification and partial identification of principal causal effects, particularly in relation to the principal ignorability assumption. We focus on the aforementioned model-based approaches that jointly model the potential intermediates and observed outcome as these approaches are commonly used, and as we will show, have inherent difficulties with respect to identification that must be managed carefully. Our first contribution is to discuss the role of principal ignorability in estimating the unknown correlation parameter between the two potential intermediate variables. Recent works have estimated this parameter by jointly modeling the potential intermediates and observed outcome, and we study this process theoretically under a Bayesian model. We show that the posterior distribution of this crucial, and typically unidentified, parameter is indeed consistent for the true correlation parameter, but only if the outcome model is known and principal ignorability is violated. We derive the asymptotic variance of this posterior distribution providing insights on when this parameter can be estimated efficiently. While this is a promising result, extending to the case where the outcome model is estimated is more challenging, as our subsequent results show that even when the correlation parameter is known, an outcome model that allows the outcome to depend on both potential intermediate variables is necessarily only partially identifiable, even under extremely restrictive parametric assumptions. We discuss alternative assumptions that can be made to make these partial identification regions more informative, or even obtain point identification. Lastly, we extend these results to more flexible semiparametric and nonparametric Bayesian models.
2 Notation and review of existing work
2.1 Notation and estimands
Throughout let $\boldsymbol{X}$ denote a $p$-dimensional vector of pre-treatment covariates, $T$ a binary treatment, $S$ a binary or continuous intermediate, and $Y$ an outcome of interest. Our observed data consists of $n$ independent copies of these random variables, and we refer to the observed data by $\mathcal{D}$. We adopt the potential outcome framework [34] and define $S(t)$ to be the potential intermediate that would be observed had treatment been set to $t$, and similarly let $Y(t)$ be the potential outcome. Throughout, we let $\boldsymbol{U}=(S(0),S(1))$ denote the joint values of the two potential intermediates, and let $S$ be the observed value of the intermediate. We assume the Stable Unit Treatment Value Assumption (SUTVA) [35], which implies that 1) there is no ’interference,’ meaning that the potential outcomes for unit $i$ do not change with the treatments assigned to other units, and 2) there are not different versions of each treatment level. This assumption ensures that $S=S(T)$ and $Y=Y(T)$. Principal strata are defined by the joint potential values of the intermediate variable $\boldsymbol{U}=(S(0),S(1))$. The causal estimands of interest are then defined in terms of average differences in potential outcomes within principal strata, which are given by
$PCE(\boldsymbol{u})=E[Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u}].$
One may also be interested in these effects conditional on covariates, which are given by
$PCE(\boldsymbol{u},\boldsymbol{x})=E[Y(1)-Y(0)\mid\boldsymbol{U}=\boldsymbol{u},\boldsymbol{X}=\boldsymbol{x}].$
Now that these effects are defined, we can review standard identification assumptions used to allow these to be identified from the observed data.
2.2 Review of identification assumptions in PS analysis
A crucial assumption for identifying causal effects is that the treatment is exchangeable given covariates, or that there are no unmeasured confounders, which is given in the following assumption.
Assumption 1 (Treatment Ignorability).
$T\perp\!\!\!\!\perp(S(0),S(1),Y(0),Y(1))\mid\boldsymbol{X}$.
Treatment ignorability is a standard assumption and it holds by design in a randomized experiment. Importantly, it ensures that there are no unmeasured confounders of the treatment-intermediate and treatment-outcome relationship. Additionally, we must make a positivity assumption ensuring that treatment can take either value for any unit in our population.
Assumption 2 (Positivity).
0 ¡ $P(T=1\mid\boldsymbol{X}=\boldsymbol{x})<1$ for all $\boldsymbol{x}$.
These two assumptions, combined with the SUTVA assumption, allow for identification of average causal effects, but they are generally not sufficient for identification of principal causal effects. In what follows, we detail different identification assumptions for PCEs separated by the binary and continuous intermediate settings, where these assumptions differ.
2.2.1 Binary intermediates
Common assumptions with binary intermediates are monotonicity, exclusion restrictions, and principal ignorability.
Assumption 3 (Monotonicity).
$S(1)\geq S(0)$.
The monotonicity assumption, sometimes referred to as the no defiers assumption, is plausible in many applications. For instance, in studies of noncompliance, treatment is not available to those who are randomly assigned to the control condition. This assumption rules out the principal strata ($S(0)=1,S(1)=0$), which leads to identification of the distribution of principal strata. This assumption is commonly used along with the following assumption:
Assumption 4 (Exclusion Restriction).
There are two exclusion restrictions:
  1. 1.
    Exclusion restriction for units of type 00: $Y(0)=Y(1)$ if $\boldsymbol{U}=(0,0)^{T}$.
  2. 2.
    Exclusion restriction for units of type 11: $Y(0)=Y(1)$ if $\boldsymbol{U}=(1,1)^{T}$.
This assumption rules out a direct effect of the treatment on the outcome that does not go through a change in the intermediate. Again, this is reasonable in randomized studies of noncompliance where randomization to a treatment should not affect the outcome, except through its impact on what treatment is actually taken. Under the exclusion restriction and monotonicity assumptions, the average causal effect for compilers is identified as the ratio of the average causal effects on $Y$ and $S$:
$E[Y(1)-Y(0)\mid S(0)=0,S(1)=1]=\frac{E[Y(1)-Y(0)]}{E[S(1)-S(0)]}.$
While plausible in certain settings, such as those with randomized treatment assignment and noncompliance, these are frequently not likely to hold, as there are many settings where a direct effect of treatment is plausible. In such settings, one can instead use the following assumption
Assumption 5 (Principal Ignorability).
$E[Y(t)\mid S(t),S(1-t),\boldsymbol{X}]=E[Y(t)\mid S(t),\boldsymbol{X}],\qquad t=0,1.$
Note that researchers sometimes invoke a slightly stronger assumption of full conditional independence, though conditional mean independence is sufficient for all estimands considered here so we proceed in this manner. One implication of this assumption is that $\boldsymbol{X}$ includes all confounders between the intermediate and outcome. Alternatively, the principal ignorability assumption can be viewed as a homogeneity assumption, as it implies that the potential outcome means are the same across certain principal strata. This assumption greatly facilitates identification because under principal ignorability and treatment ignorability, we have that the PCE for principal strata $\boldsymbol{u}=(s_{0},s_{1})$ is identified as
$PCE(\boldsymbol{u},\boldsymbol{x})=E[Y\mid S=s_{1},\boldsymbol{X}=\boldsymbol{x},T=1]-E[Y\mid S=s_{0},\boldsymbol{X}=\boldsymbol{x},T=0],$
which is a function of observable variables. The identification of the average PCE requires integration of this quantity over the covariate distribution, which requires knowledge of the conditional probabilities of principal strata, $P(\boldsymbol{U}\mid\boldsymbol{X})$, also known as the principal score. The identification of the principal score requires additional assumptions beyond principal ignorability.
2.2.2 Continuous intermediates
For a binary $S$, there are only four principal strata, and the monotonicity assumption eliminates one of these groups and enables the identification of the probability mass of $\boldsymbol{U}\mid\boldsymbol{X}$ ([1]). For a continuous $S$, there are infinitely many principal strata, and while the monotonicity assumption can rule out some of them, there still remain infinitely many principal strata. Thus, monotonicity alone cannot produce the identification of the distribution of $\boldsymbol{U}\mid\boldsymbol{X}$, sometimes referred to as the principal density for continuous intermediates. Typically, researchers make parametric assumptions in order to identify this distribution. One approach is to assume the joint distribution of principal strata follows a known copula function $\mathbb{C}_{\rho}(\cdot,\cdot)$ (see [28]; [20]; [4])
$P(S(0)\leq s_{0},S(1)\leq s_{1}\mid\boldsymbol{X})=\mathbb{C}_{\rho}(P(S(0)\leq s_{0}\mid\boldsymbol{X}),P(S(1)\leq s_{1}\mid\boldsymbol{X})),$
where the parameter $\rho$ indicates the correlation between $S(0)$ and $S(1)$ given $\boldsymbol{X}$.
In general, the explicit identification of the correlation between potential intermediates is infeasible because only one of them is ever observed per observation. This has led to a common practice of $\rho$ being selected based on domain knowledge, and a sensitivity analysis is performed to assess the impact of this parameter. Throughout this article, we refer to parameters like $\rho$ that cannot be identified marginally as the association parameters of the principal strata. Under this distributional assumption, combined with Assumptions (1) and (5), [28] demonstrated nonparametric identification of PCEs. Given that knowledge of this correlation parameter is a strong assumption, and that results can be sensitive to its choice, many authors have taken a parametric modeling approach to estimate this correlation. Additionally, in most of these works, principal ignorability is not assumed, showing that the parametric modeling assumptions reduce the number of structural assumptions that need to be made. This is the focus of our work, as we describe the extent to which inference in these settings is possible in the following sections.
3 On the possibility of identifying $\rho$
3.1 A simple Bayesian parametric model
Most approaches that jointly model the observed outcome and potential intermediates are Bayesian, which will be the focus of our manuscript, however, the results have similar implications for frequentist approaches aiming to estimate $\rho$, such as those seen in [4]. Bayesian approaches are popular for this choice ([21], [37], [23], [2]) as they naturally handle missing data and can propagate uncertainty from missing data imputations into the final causal estimands. This general approach consists of two parts: one is the specification of a model for the observed outcome and the other is the specification of a model for the joint values of the two potential intermediates, i.e., the principal strata.
For simplicity of discussion and derivations, we first focus on a simple, linear parametric model for the outcome as well as the principal strata. We extend these ideas to semiparametric and nonparametric models in Section 7. We focus on a model given by
$\begin{split}E[Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]&=\boldsymbol{\beta}_{t}^{T}\boldsymbol{U}_{i}+\lambda_{t}+\boldsymbol{\gamma}^{T}\boldsymbol{X}_{i},\\ E[\boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}]&=(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i})^{T},\end{split}$ (1)
where $\boldsymbol{\beta}_{t}=(\beta_{t0},\beta_{t1})$ for $t=0$, 1. It may appear that we are focusing on overly simple models, but in what follows, we will see that even in such simple, parametric situations, issues of model identifiability arise, which will have implications for more complex model structures.
As mentioned earlier, we primarily focus on settings with a continuous intermediate variable, though analogous results hold for the binary intermediate variable setting. We assume Gaussian errors in the outcome model and a joint Gaussian distribution for the principal strata. We define $\mu_{yit}$ to be the conditional mean of the potential outcomes, $E[Y_{it}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]$, and $\boldsymbol{\mu}_{si}=(\mu_{si0},\mu_{si1})$ to be the conditional mean of principal strata. Thus, a full data generating model can be written as:
$\begin{split}Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t&\sim N(\mu_{yit},\sigma_{y}^{2}),\qquad t=0,1.\\ \boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}&\sim N(\boldsymbol{\mu}_{si},\boldsymbol{\Sigma}_{s}),\end{split}$ (2)
where
$\boldsymbol{\Sigma}_{s}=\begin{pmatrix}\sigma_{s0}^{2}&\rho\sigma_{s0}\sigma_{s1}\\ \rho\sigma_{s0}\sigma_{s1}&\sigma_{s1}^{2}\end{pmatrix}.$
In model (2), $\rho$ is the association parameter of the principal strata and plays an important role in estimating PCEs. To the best of our knowledge, there is no literature formally studying whether it is possible to identify the association parameters, which we do in the following section.
3.2 Identification of the association parameter of the principal strata
For a continuous intermediate, estimating average PCEs always requires estimating the principal density, which involves the association parameters. Identifying the association parameters is generally infeasible due to the lack of joint observations of principal strata. Some researchers assume these are known ([21], while others have tried to estimate them using similar models to those in 2 ([37], [4]). The idea is that although $S(0)$ and $S(1)$ are never jointly observed, their association is implicitly included in the outcome model. The following result provides theoretical justification to this strategy:
Theorem 1.
Suppose that $\boldsymbol{\theta}=(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{1},\lambda_{0},\lambda_{1},\sigma_{y}^{2},\boldsymbol{\gamma},\boldsymbol{\alpha},\sigma_{s0},\sigma_{s1})$ are known and principal ignorability fails: that is, at least one of $\beta_{01}$ or $\beta_{10}$ is nonzero. Under model 2, the posterior mode of $\rho\mid\mathcal{D}$ is consistent.
Theorem 1 shows when principal ignorability is violated, then identification of $\rho$ is indeed possible, though this result should be interpreted with caution. For one, it required the outcome model to be correctly specified with known parameters. We will see in the subsequent sections that the model parameters are only partially identified in many settings, and the implications of Theorem 1 would only be useful for estimating $\rho$ if we can consistently estimate these regression parameters. Another issue is that this result relies on principal ignorability being violated. One can show that the posterior of $\rho$ reduces to the prior distribution for $\rho$ when principal ignorability holds. This can also be seen in Theorem 2, where we show the asymptotic variance of the posterior distribution of $\rho$.
Theorem 2 (Asymptotic approximation of posterior variance).
Suppose that $\boldsymbol{\theta}$ are known and principal ignorability fails: that is, at least one of $\beta_{01}$ or $\beta_{10}$ is nonzero. Under model (2), an asymptotic approximation of the posterior variance of $\rho\mid\mathcal{D}$ is
$\begin{split}Var(\rho\mid\mathcal{D})&\approx n^{-1}\bigg{(}\bar{T}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}\bigg{(}\frac{2{{\rho}^{\ast}}^{2}{\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}}{(\zeta_{1}^{*}-{\psi_{1}^{*}}^{2})^{2}}+\frac{1}{\zeta_{1}^{*}-{\psi_{1}^{*}}^{2}}\bigg{)}+\\ &\qquad(1-\bar{T}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}\bigg{(}\frac{2{{\rho}^{\ast}}^{2}{\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}}{(\zeta_{0}^{*}-{\psi_{0}^{*}}^{2})^{2}}+\frac{1}{\zeta_{0}^{*}-{\psi_{0}^{*}}^{2}}\bigg{)}\bigg{)}^{-1},\end{split}$
where $\bar{T}=\sum\limits_{i=1}^{n}T_{i}/n$, $\zeta^{*}_{0}-{\psi^{*}_{0}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{01}^{\ast}}^{2}{\sigma_{s1}^{\ast}}^{2}$, and $\zeta_{1}^{*}-{\psi^{*}_{1}}^{2}={{\sigma}_{y}^{\ast}}^{2}+(1-{{\rho}^{\ast}}^{2}){\beta_{10}^{\ast}}^{2}{\sigma_{s0}^{\ast}}^{2}$.
Note that asterisks represent the true parameter values. We can see that ${\beta_{01}^{\ast}}$ and ${\beta_{10}^{\ast}}$ appear in the denominator of the approximation for the posterior variance. These two parameters together indicate the level of violations of principal ignorability, and we see that when both these parameters are close to zero, the posterior variance will be very large. Theorems 1 and 2 together show that either principal ignorability holding, or even very slight violations of principal ignorability, will not be sufficient for accurately estimating the unknown correlation $\rho$. Overall, these results show that identification of $\rho$ is possible, but only if the outcome model is correctly specified and its parameters are able to be consistently estimated, and if principal ignorability fails. In the following section, we study this first issue of identifying the parameters of the outcome model in more detail when principal ignorability is violated.
4 Partial identification when principal ignorability fails
4.1 Partial identification and principal causal effects
Partial identification commonly arises in causal inference problems due to the inherent missingness of potential outcomes. Identification can commonly be achieved under stronger assumptions, but these are not always plausible and incorrectly assuming them can lead to bias when estimating causal effects. In Section 2, we reviewed common assumptions for binary and continuous intermediate variables that can lead to identification of principal causal effects. Within the context of continuous intermediates, nonparametric identification has been established under both a principal ignorability assumption, as well as a distributional assumption on the potential intermediate that assumes $\rho$ is known [28, 42]. Other papers have estimated principal causal effects in this setting without assuming principal ignorability [23, 25], and at times also not assuming knowledge of $\rho$ [37, 4]. In this section, we bridge this gap by providing information about what is possible when principal ignorability is not assumed, and whether then $\rho$ can be subsequently estimated as suggested by the existing literature and the results in Section 3. We do so in the context of the simple, linear model presented in (2), but as we see in Section 7, these results have implications for a much broader class of models.
We now show that the parameters in Model (2), which does not assume principal ignorability, are generally not identifiable even in this simple, parametric setting when $\rho$ is also treated as known. Of particular interest are the parameters dictating violations of principal ignorability, given by $(\beta_{01},\beta_{10})$. To see this, we can look at the distribution of the observed data, conditional on $T_{i}=t$ and $\boldsymbol{X}_{i}=\boldsymbol{x}$, which is given by
$(Y_{i},S_{i})\mid T_{i}=t,\boldsymbol{X}_{i}=\boldsymbol{x}\sim N\Bigg{(}(\mu_{yt},\mu_{sit}),\begin{pmatrix}\zeta_{t}&\psi_{t}\sigma_{st}\\ \psi_{t}\sigma_{st}&\sigma^{2}_{st},\end{pmatrix}\Bigg{)},$ (3)
where $\mu_{yt}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0}+\boldsymbol{\alpha}^{T}\boldsymbol{x},\phi_{1}+\boldsymbol{\alpha}^{T}\boldsymbol{x})+\boldsymbol{\gamma}^{T}\boldsymbol{x}$, $\mu_{sit}=\phi_{t}+\boldsymbol{\alpha}^{T}\boldsymbol{x}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$ for $t$ = 0, 1. Clearly the parameters of the observed data distribution are identifiable from the observed data, and given $\boldsymbol{X}_{i}=\boldsymbol{x}$, there are 5 identifiable parameters for each of $t=0,1$ leading to 10 identifiable parameters. Excluding the parameters corresponding to $\boldsymbol{X}$ given by $(\boldsymbol{\gamma},\boldsymbol{\alpha})$, Model (2) has 11 parameters, showing that there are fewer known equations than unknown variables. This leads to partial identification of certain parameters, despite being the simplest, parametric model possible that allows for violations of principal ignorability. Partial identification means that there are multiple points, or a region of parameter values, that are equally supported by the data and lead to the same observed data distribution, which can not be distinguished even with an infinite sample size. For a general discussion on Bayesian inference in such scenarios, we point readers to [12]. Naturally, this also implies that the parameters of more complex semi- or nonparametric models, such as those seen in the literature [4, 37], are at best partially identified in this setting unless additional constraints are imposed on the model parameters.
In the presence of such partial identification, we are left with two options. One can proceed with Bayesian inference where the parameters are only partially identified, and we detail the size of such partial identification regions in Section 4.2. Alternatively, additional assumptions can be made, which allows for point identification of the model parameters, and potentially $\rho$ as well, which we detail in Section 4.3. We also provide weaker, alternative assumptions in Section 5 that do not obtain point identification, but can drastically reduce the widths of partial identification regions.
4.2 Partial identification regions when principal ignorability fails
Throughout the rest of this paper, we focus on partial identification regions for parameters of interest without considering prior distributions for these parameters. Partial identification regions are equally supported by the observed data, but the posterior is not necessarily flat in these regions, due to the influence of the prior distribution. Clearly, the prior distribution affects the posterior distribution, and if strong prior knowledge is available, then this should be incorporated. We assume that such knowledge is not available, and that flat priors are placed on unidentified parameters. Using $(\beta_{01},\beta_{10})$ and the identifiable parameters in (3), the PCEs of interest can be written as
$\begin{split}PCE(\boldsymbol{u})&=(\beta_{10}-\frac{\psi_{0}}{\sigma_{s0}}+\frac{\sigma_{s1}}{\sigma_{s0}}\rho\beta_{01})(s_{0}-\phi_{0})+\\ &\qquad(\frac{\psi_{1}}{\sigma_{s1}}-\frac{\sigma_{s0}}{\sigma_{s1}}\rho\beta_{10}-\beta_{01})(s_{1}-\phi_{1})+(\mu^{\prime}_{y1}-\mu^{\prime}_{y0}),\end{split}$ (4)
where $\mu^{\prime}_{yt}=\lambda_{t}+\boldsymbol{\beta}_{t}^{T}(\phi_{0},\phi_{1})$ for $t=0,1$, and $(\beta_{01},\beta_{10})$ are subject to the constraints imposed by (2) and (3). Note that $\mu^{\prime}_{yt}$ is identifiable and corresponds to the mean of potential outcomes, excluding the effects of covariates. Details about their identification are provided in Appendix A. The PCEs depend on the principal strata $\boldsymbol{u}=(s_{0},s_{1})$ and on $(\beta_{01},\beta_{10})$, and therefore we aim to study the partial identification region of $(\beta_{01},\beta_{10})$. For simplicity of the expressions for the partial identification regions of $(\beta_{01},\beta_{10})$, we exclude covariates from the theoretical results regarding partial identification regions. Note that in the appendix we show that to incorporate covariates, we only need to condition additionally on $\boldsymbol{X}$ for all conditional variances or correlations found in the partial identification regions
Proposition 1.
Suppose that $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$. Then the partial identification region of $\beta_{01}$ is
$\left[-\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}},\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\right],$
and the partial identification region of $\beta_{10}$ is
$\begin{split}&\left[-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}},-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}}\right]\cup\\ &\qquad\left[\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}}\right],\end{split}$
where $\beta_{01}$ and $\beta_{10}$ also satisfy the constraint
$\text{Var}(S(0))\beta_{10}^{2}-\text{Var}(S(1))\beta_{01}^{2}=\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{1-\rho^{2}}.$ (5)
Note that if $\text{Var}(Y(0)\mid S(0))\geq\text{Var}(Y(1)\mid S(1))$, one can simply exchange 0 and 1 in Proposition 1 to obtain the corresponding partial identification region. Also note that all terms in the partial identification region are identifiable since we are assuming for now that $\rho$ is known. Also, an interesting consequence of equation (5) is that if the right-hand side is non-zero, then at least one of $\beta_{01}$ and $\beta_{10}$ must be nonzero, indicating principal ignorability is necessarily violated. Proposition 1 shows that both the sign and magnitude of ($\beta_{01},\beta_{10}$) are not identified, and that these regions can be rather large, which will lead to large partial identification regions for the PCEs of interest. Further, we can see an interplay between the size of the partial identification region for the two parameters. When the partial identification region for $\beta_{01}$ is larger due to a large value of $\text{Var}(Y(0)\mid S(0))$, this leads to a reduction in the width of the region for $\beta_{10}$, and vice-versa. This interplay also effectively guarantees that the partial identification region for at least one of the parameters (if not both) is large, and therefore principal causal effects will have wide partial identification regions as well.
4.3 Assumptions for identification when principal ignorability is violated
The previous sections cast doubt on the ability to perform inference when principal ignorability is violated, and additionally question whether the results from Section 3.2 are useful in practice, because they showed that $\rho$ can only be identified under known outcome model parameters. Here, we discuss different assumptions that can not only lead to the identification of PCEs given known $\rho$, but also to the identification of $\rho$ itself. As discussed earlier, there are 10 identified parameters in (3), while Model (2) has 11 unknown parameters if $\rho$ is known, or 12 unknown parameters otherwise. The easiest way to make these parameters point identifiable is to add constraints on the parameters of Model (2) that reduce the number of unknown parameters. For instance, principal ignorability can be viewed as imposing the constraint $\beta_{01}=\beta_{10}=0$, which leads to the identification of all parameters in Model (2). Despite there being two constraints imposed, this particular assumption does not identify $\rho$, because Theorem 1 showed that principal ignorability must be violated to obtain identification of $\rho$.
Alternative assumptions beyond the principal ignorability assumption have been proposed in the literature, which we can adopt to obtain point identification. One such modeling assumption explored in recent work [37, 4] in our setting amounts to setting $\beta_{00}=\beta_{10},\beta_{01}=0$. This assumption treats $E[Y(0)\mid\boldsymbol{U},\boldsymbol{X}]$ as a baseline characteristic that is shared between the treatment and control groups [37]. This assumption identifies $\rho$ only up to its sign. However, it is typically reasonable to assume $\rho>0$, which enables full identification of $\rho$. Alternatively, one could make the assumption that both $\lambda_{0}=\lambda_{1}$ and $\beta_{01}=0$, which amounts to assuming there is no direct effect of treatment on the outcome and that the unobserved intermediate does not affect the outcome in the control group. Note that both of these possibilities impose two constraints on the parameter space, which allows us to identify both the outcome model parameters and the unknown correlation $\rho$. If $\rho$ is treated as known, or as a sensitivity parameter to be varied, then only one constraint is needed to obtain identification of the outcome model parameters. It is important to emphasize that while these constraints can be placed to obtain identification, they should only be made if they are deemed plausible in the application to which they are utilized. If one can not make such strong assumptions, then one should continue with partial identification. To this end, we now shift focus to incorporating more plausible assumptions that do not obtain point identification, but can reduce the widths of partial identification regions significantly when such strong identifying assumptions are not plausible.
5 Weaker alternative assumptions of principal ignorability
Strong assumptions like principal ignorability or the assumptions discussed in Section 4.3 can lead to point identification of principal causal effects, but they are untestable and may be overly restrictive in many settings. Without them, however, partial identification regions such as those seen in Proposition 1 can be overly wide and uninformative even in contexts with simple, parametric models. This is expected to persist or be exacerbated in the context of more flexible outcome models. To allow for violations of principal ignorability and reduce the widths of partial identification regions, we propose two weaker assumptions that can be reasoned about in any particular application. The first assumption addresses the issue of unidentified signs in the effects of the unobserved intermediate.
Assumption 6 (Same sign).
Assume
$\text{sign}\Bigg{(}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S(T)}\Bigg{)}=\text{sign}\Bigg{(}\dfrac{\partial E(Y\mid T,\boldsymbol{U},\boldsymbol{X})}{\partial S_{1-T}}\Bigg{)}.$
Assumption 6 states that, given the covariates and treatment, the observed and unobserved intermediates should have effects in the same direction on the outcome. This is plausible in many applications where the effect of the intermediate on the outcome can only plausibly be in one direction. In the case of Model (2), this simplifies to assuming that $\beta_{01}$ and $\beta_{00}$ have the same sign, and that $\beta_{10}$ and $\beta_{11}$ have the same sign. Clearly this assumption rules out a significant portion of the partial identification region found in Proposition 1, which is shown in the following result. Under Assumption 6, only the magnitudes of ($\beta_{01},\beta_{10}$) are unidentified.
Proposition 2.
Suppose that Assumption 6 holds and $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$. Assume $\beta_{00}\geq 0$ and $\beta_{11}\geq 0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}0,\sqrt{\frac{\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\Bigg{]},$
and the partial identification region for $\beta_{10}$ is
$\Bigg{[}\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]},$
where $\beta_{01}$ and $\beta_{10}$ also satisfy (5).
For $\beta_{00}$ and $\beta_{11}$ with signs different from those assumed in Proposition 2, the corresponding partial identification regions for $\beta_{01}$ and $\beta_{10}$ can be obtained by flipping the signs to match those of $\beta_{00}$ and $\beta_{11}$, respectively. This result also requires us to know the sign of $\beta_{tt}$ for $t=0,1$. If this is not known a priori, then one can regress the observed $Y$ on the observed $S$ and $\boldsymbol{X}$, given $T=t$, to determine the sign for $\beta_{tt}$. This approach is not guaranteed to correctly identify the sign, but will only fail to do so in certain extreme situations, such as when $\rho$ is close to $-1$. This cuts the partial identification region in half, providing large benefits in the uncertainty in the resulting causal estimates. While these benefits may be large, it is not possible in many situations to know this assumption is true a priori. Now we present a second, weaker assumption, which should hold true in most applications. This second assumption effectively assumes that the unobserved intermediate, once we condition on the observed one, has a smaller impact on the outcome than the observed one had. This can be seen as a weakened version of principal ignorability as the principal ignorability assumption assumes that there is no effect of the unobserved intermediate, once we condition on the observed one. This is formalized in the following assumption.
Assumption 7 (Dominant Observed Effect).
Assume
$R^{2}_{Y\sim S(1-T)\mid S(T),T,\boldsymbol{X}}\leq R^{2}_{Y\sim S(T)\mid T,\boldsymbol{X}}.$
Assumption 7 implies that the signal-to-noise ratio (SNR) of the unobserved intermediate, conditional on the observed intermediate, is no greater than that of the observed intermediate. This assumption relaxes principal ignorability by allowing the unobserved intermediate to affect the outcome, but precludes large effects of the unobserved intermediate. Although Assumption 7 is formulated in terms of Pearson’s $R^{2}$, one could also formulate the assumption in terms of nonparametric partial $R^{2}$ values that would apply in broader contexts, as we see in Section 7. This assumption reduces the width of the partial identification region substantially as shown in the following result.
Proposition 3.
Suppose that Assumption 7 holds, $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$, $\frac{(\text{Var}(Y(0)\mid S(0)))^{2}}{\text{Var}(Y(0))}\geq\frac{(\text{Var}(Y(1)\mid S(1)))^{2}}{\text{Var}(Y(1))}$, and $\beta_{00}>0$. The partial identification region for $\beta_{01}$ is
$\Bigg{[}-\sqrt{\frac{\text{Var}(Y(0)\mid S(0))\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}(S(1))}},\sqrt{\frac{\text{Var}(Y(0)\mid S(0))\text{Cor}^{2}(Y(0),S(0))}{(1-\rho^{2})\text{Var}(S(1))}}\Bigg{]},$
and the partial identification region for $\beta_{10}$ is
$\begin{split}&\Bigg{[}-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}^{2}(Y(0)\mid S(0))/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}(S(0))}},-\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]}\cup\\ &\qquad\Bigg{[}\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}(Y(0)\mid S(0))}{(1-\rho^{2})\text{Var}(S(0))}},\sqrt{\frac{\text{Var}(Y(1)\mid S(1))-\text{Var}^{2}(Y(0)\mid S(0))/\text{Var}(Y(0))}{(1-\rho^{2})\text{Var}(S(0))}}\Bigg{]},\end{split}$
where $\beta_{01}$ and $\beta_{10}$ also satisfy (5).
Note that the result depends on $\text{Var}(Y(1)\mid S(1))\geq\text{Var}(Y(0)\mid S(0))$, ${(\text{Var}(Y(0)\mid S(0)))^{2}}\break/{\text{Var}(Y(0))}\geq{(\text{Var}(Y(1)\mid S(1)))^{2}}/{\text{Var}(Y(1))}$, and $\beta_{00}>0$ all being true. We should emphasize that similar bounds can be derived under any combination of directions for these inequalities, but we focus on one such choice for simplicity. By comparing the results in Propositions 1 and 3, it is clear that Assumption 7 reduces the length of the partial identification region substantially. For $\beta_{01}$ the size of the region is reduced by a factor of $|\text{Cor}(Y(0),S(0))|$, which reflects the strength of the association between the outcome and the observed intermediate within the control $(T=0)$ group. When $|\text{Cor}(Y(0),S(0))|$ is very small, the benefits brought by Assumption 7 are large since it prevents assigning overly large values to the effect of the missing intermediate.
One interesting aspect of this result, which we show in the proof of Proposition 3 in the Appendix, is that Assumption 7 is equivalent to assuming
$\dfrac{(\text{Var}[Y\mid S,\boldsymbol{X},T=t])^{2}}{\text{Var}(Y\mid\boldsymbol{X},T=t)}\leq\text{Var}(Y(t)\mid S(t),S(1-t),\boldsymbol{X},T=t),$
which shows that the assumption is equivalent to a lower bound on the residual variance, $\sigma_{y}^{2}$. One important implication of this is the ability to apply such a constraint within our Bayesian modeling framework. One can first estimate this lower bound using the observed data, and can then place a truncated prior distribution on $\sigma_{y}^{2}$ that guarantees the bound holds and therefore Assumption 7 also holds. This is also important once semiparametric or nonparametric outcome models are used, because it can be more difficult in those settings to enforce restrictions on the estimated regression functions, but a truncated prior on the residual variance is similarly straightforward in that setting.
6 Simulations and applications
Here, we examine the theoretical results obtained in the previous sections as well as the practical performance of our weakened identification assumptions for estimating principal causal effects. In the first set of simulation studies, we treat $\rho$ as known, and focus on partial identification of principal causal effects of interest under varying assumptions. Then, we explore the identification of $\rho$ under varying assumptions and sample sizes. Throughout, we do not include covariates $\boldsymbol{X}$, but similar results would be obtained if covariates were additionally adjusted for. Under both scenarios, Gibbs sampling is used, with a single MCMC chain run for a total of 25,000 iterations, where the first 5,000 are burn-in, and a thinning interval of 30 is applied. The details of the Gibbs sampler are included in Appendix B.
6.1 Performance under a known $\rho$
In this scenario, the data-generating model is based on the analysis of [4] and the dataset from the Lipid Research Clinics Coronary Primary Prevention Trial (LRC-CPPT) ([10]). The parameter values are based on the estimates from previous work with small adjustments in order to match Model (2). Specifically, the data are generated from the following data generating process:
$\begin{split}Y_{i}(0)\mid\boldsymbol{U}_{i}&\sim N(-0.5+11.5S(0),14^{2}),\\ Y_{i}(1)\mid\boldsymbol{U}_{i}&\sim N(-0.5+11.5S(0)+96S(1),14^{2}),\\ \boldsymbol{U}_{i}&\sim N((0.89,0.70)^{T},\boldsymbol{\Sigma}_{S}),\end{split}$ (6)
where
$\boldsymbol{\Sigma}_{S}=\begin{pmatrix}0.25^{2}&0.75*0.25*0.25\\ 0.75*0.25*0.25&0.25^{2}\end{pmatrix}.$
The parameter values here are based on previous studies ([4], [37]), both of which conducted analyses using the same outcome models. We assume that $\rho$ is known and focus on partial identification of the PCEs under violations of principal ignorability. One can verify that Model (6) satisfies Assumption 7, but only partially satisfies Assumption 6 because $\beta_{01}=0$ and therefore does not have the same sign as $\beta_{00}$, which is positive. We generate 500 simulated datasets from Model (6) with $n=300$. For each simulated dataset, we run three MCMC chains: the first is without either Assumption 6 or 7, the second incorporates Assumption 7, and the third incorporates Assumption 6. Note that we only partially apply Assumption 6 by applying it to the signs of $(\beta_{10},\beta_{11})$, which are equal. For the model without constraints or assumptions, we consider noninformative conjugate priors as follows:
$\begin{split}&\boldsymbol{\beta}_{t}\sim N(0,10^{5}\boldsymbol{I}_{2}),\ \lambda_{t}\sim N(0,10^{5}),\ \phi_{t}\sim N(0,10^{5}),\\ &\sigma_{y}^{2}\sim IG(10^{-3},10^{-3}),\ \sigma_{s}^{2}\sim IG(10^{-3},10^{-3}),\end{split}$
where $IG(a,b)$ represents the inverse gamma distribution and $\boldsymbol{I}_{2}$ is a $2\times 2$ identity matrix. We have noticed empirically that partial identification can lead to inefficient MCMC sampling with parameters getting trapped at extreme values. In this case, the residual variance $\sigma_{y}^{2}$ becomes trapped at values close to 0, so a lower bound of $0.05\min_{t=0,1}\text{Var}(Y\mid T=t)$ is used throughout when updating $\sigma_{y}^{2}$. This improves the poor mixing, but the issue still persists to a lesser degree. For MCMC chains under Assumption 7, since Assumption 7 is equivalent to adding a lower bound for $\sigma_{y}^{2}$, then a truncated inverse gamma distribution for $\sigma_{y}^{2}$ is utilized, where a lower bound for this truncation is the empirical estimate of $0.9\min_{t=0,1}\left((\text{Var}[Y\mid S,T=t])^{2}/\text{Var}(Y\mid T=t)\right)$. For MCMC chains that partially incorporate Assumption 6, a truncated normal prior for $\boldsymbol{\beta}_{1}$ is used to ensure that $(\beta_{10}>0,\beta_{11}>0)$.
We target PCEs defined on ($S(0),S(1)$) at the combinations of the quartiles of the intermediate values, as used by [21, 4, 37]. Since the true values of the PCEs in this scenario depend only on $S(1)$, we fix $S(0)$ at its median value. The simulation results are shown in Table 1. All empirical coverage rates (ECRs) are above 95%, indicating that valid credible intervals (CIs) are obtained under any of the proposed assumptions. The average widths of CIs were compared, and the proposed assumptions significantly reduced average widths. Particularly, for $S(1)$ at the first and third quartiles, Assumption 7 provided 42% reductions in the average width and Assumption 6 provided 26% reductions relative to not making either assumption.
No constraints Dominant effect Same sign
True value Mean ECR Width Mean ECR Width Mean ECR Width
$\boldsymbol{u}=(0.89,0.18)$ 17 17 0.99 43 16 1.00 25 23 0.97 32
$\boldsymbol{u}=(0.89,0.35)$ 34 34 0.99 13 34 1.00 9 34 0.98 12
$\boldsymbol{u}=(0.89,0.52)$ 50 50 0.99 43 51 1.00 25 45 0.98 32
Table 1: Comparisons of average posterior mean (Mean), empirical coverage rates (ECR) and average width of credible intervals (Width) under varying assumptions that allow for violations of principal ignorability.
6.2 Identifying $\rho$ under different constraints
Now we use a related data generating process to further explore the ability to identify $\rho$ under varying assumptions. As discussed previously, one plausible identification assumption for $\rho$ is $\beta_{01}=0$ and $\beta_{00}=\beta_{10}$. This guarantees the identification of $\rho$ when we also assume $\rho>0$. To investigate both identification and posterior consistency, we vary $n$ within {300, 600, 1200} and we generate one dataset for each sample size. For each dataset, we run MCMC under no constraints, one constraint ($\beta_{01}=0$), and two constraints ($\beta_{01}=0$, $\beta_{00}=\beta_{10}$). To avoid the sign issue, we also constrain $\rho$ to be in $(0.00,0.95)$, where setting $\rho<0.95$ prevents MCMC sampling from getting trapped around the extreme value $\rho=1$.
We generate data from the following data generation process:
$\begin{split}Y_{i}(0)\mid\boldsymbol{U}_{i}&\sim N(0.9+1.2S(0),0.5^{2}),\\ Y_{i}(1)\mid\boldsymbol{U}_{i}&\sim N(0.5+1.2S(0)+1.2S(1),0.5^{2}),\\ \boldsymbol{U}_{i}&\sim N((0.3,0.5)^{T},\boldsymbol{\Sigma}_{S}),\end{split}$ (7)
where $\boldsymbol{\Sigma}_{S}=\begin{pmatrix}1&0.75\\ 0.75&1\end{pmatrix}$. The same noninformative conjugate priors are used and we additionally place a flat prior on the correlation between potential intermediates, $\rho\sim U(0,0.95)$.
The posterior distributions of $\rho$ across all three scenarios and sample sizes are shown in Figure 1. Regardless of sample size, there is a notable spike around the true value of $\rho=0.75$ when we place two constraints on the model parameters. There appear to be mild spikes close to 1 when there are no constraints or one constraint placed on the model, but this is simply due to poor mixing as the MCMC sampler gets stuck at $\rho$ values close to 1. When we alternatively assume $\rho\sim U(0,0.9)$, these spikes are reduced. Importantly, as $n$ increases, their height does not increase, whereas the height of spike under two constraints increases significantly, which points towards consistency of the correlation parameter. In Appendix C, we include more plots of the estimated posterior variance for more sample sizes, which show that asymptotically, the posterior variance is $O(n^{-1})$.
Figure 1: The posterior distributions of $\rho$ corresponding to $n=$300, 600, 1200 under three different amounts of constraints on the model parameters.
7 Extensions to Bayesian nonparametric modeling
In Sections 4 and 5, we discussed partial identification regions and alternative assumptions when principal ignorability fails. While many of these results were described in terms of conditional variances not unique to a specific model, they were derived under the assumption of a linear outcome model. This facilitated derivations and ease of exposition, but now we show that many similar ideas hold in more general modeling contexts, which has broader implications for the analysis of principal causal effects when principal ignorability is violated.
We consider directly extending Model (2) to a more flexible modeling framework that does not make the same, potentially restrictive, parametric assumptions. Specifically, we can model the outcomes as follows:
$\begin{split}Y_{i}\mid\boldsymbol{U}_{i},\boldsymbol{X_{i}},T_{i}=t&\sim\mu_{yt}(\boldsymbol{X}_{i},S_{i})+\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})+\epsilon_{i},\\ \boldsymbol{U}_{i}\mid\boldsymbol{X}_{i}&\sim N(\boldsymbol{\mu}_{s}(\boldsymbol{X}),\boldsymbol{\Sigma}_{s}),\end{split}$ (8)
where $\epsilon_{i}\sim N(0,\sigma_{y}^{2})$, $\epsilon\perp\!\!\!\!\perp\boldsymbol{U}_{i}$, and $E[Y_{it}\mid S_{i},\boldsymbol{X_{i}},T_{i}=t]=\mu_{yt}(\boldsymbol{X}_{i},S_{i})$. The subscript $c$ implies that $E[\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})\mid S_{i},\boldsymbol{X}_{i}]=0$, which means it is centered. Therefore we can view $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})$ as a nonparametric extension to $\beta_{t,1-t}S(1-t)$ as it controls the degree of violations to principal ignorability, and if $\mu_{yct}(\boldsymbol{X}_{i},\boldsymbol{U}_{i})=0$, then principal ignorability holds. Our results here are general and hold for any nonparametric Bayesian prior placed on these functions, but either Gaussian processes or Bayesian additive regression tree (BART, [7]) priors would likely be used. The latter of which has been used for estimating principal causal effects [25] and is commonly used more generally within causal inference [15, 13, 27].
7.1 Partial identification related to extended sign issues
Without principal ignorability, one aspect of partial identification concerns the unidentified signs within a parametric linear outcome model. In the nonparametric case, this concept generalizes to involve a specific class of transformations known as invariant transformations.
Definition 1.
A transformation $M$ of random variable $Z$ is said to be an invariant transformation if $M(Z)$ and $Z$ follow the same distribution.
Let $\mathcal{M}(Z)$ denote the collection of all invariant transformations of a random variable $Z$. To formalize the role of invariant transformations in identifying distributions, we present the following result, which shows that principal causal effects are not identified in this setting.
Theorem 3.
For any invertible $M\in\mathcal{M}(S(1-t)\mid S(t),\boldsymbol{X})$, let $Y^{\prime}(t)=\mu_{yt}(\boldsymbol{X},S(t))+\mu_{yct}(\boldsymbol{X},S(t),M^{-1}(S(1-t)))+\epsilon$. Then, marginally, $(Y^{\prime}(t),S(t))\mid\boldsymbol{X}$ and $(Y(t),S(t))\mid\boldsymbol{X}$ follow the same distribution.
Theorem 3 shows that $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ is unidentifiable, and therefore $E[Y(t)\mid S(t),S(1-t),\boldsymbol{X}]$ and any PCEs of interest are also unidentifiable. Because of the unobserved nature of $S(1-t)$, the flexibility of nonparameteric outcome models allows various possible ways for $S(1-t)$ to contribute to the observed outcome. These plausible outcome models can lead to completely different or even contradictory PCEs. For instance, if $(S(t),S(1-t))$ follows a bivariate Gaussian distribution, then one such invariant transformation can be constructed as $M:s\to 2m_{1-t}(s_{t},\boldsymbol{x})-s$, where $m_{1-t}(s_{t},\boldsymbol{x})$ denotes $E[S(1-t)\mid S(t)=s_{t},\boldsymbol{X}=\boldsymbol{x}]$. Clearly, $M$ is invertible and $M^{-1}=M$. Partial identification arises if $\mu_{yct}(\boldsymbol{x},s_{t},2m_{1-t}(s_{t},\boldsymbol{x})-s_{1-t})\neq\mu_{yct}(\boldsymbol{x},s_{t},s_{1-t})$, which is commonly the case. To provide intuition for this, this inequality would hold in the linear model in (2) if $\beta_{t,1-t}$ is not equal to 0, or equivalently, when principal ignorability fails.
Although we assume a bivariate Gaussian distribution for $(S(t),S(1-t))$ in Model (8), Theorem 3 does not rely on this assumption and holds for $(S(t),S(1-t))$ under any distribution. Theorem 3 generalizes the sign issues leading to partial identification in the linear model setting to a more general partial identification result based on a class of invariant transformations. Usually, the number of transformations in this class is infinite. However, most are artificial, resulting in complex and unusual potential outcome models $Y^{\prime}(t)$, which can easily be ruled out under reasonable modeling constraints. For cases like the sign issue, one can use domain knowledge to determine the correct transformation or employ reasonably uninformative priors to rule out other, incorrect outcome models.
7.2 Partial identification based on magnitude of unobserved intermediate effect
Under linear outcome models, the sign of the effect of the unobserved intermediate is not identified without additional constraints, but even if the sign were known, the magnitude of the effect is also not identified. Similarly, for nonparametric outcome models, even if the invariant transformations discussed previously were not an issue, we still are not able to identify the magnitude of the effect of the unobserved intermediate variable. In Model (8), this is represented by the magnitude of the $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$ function. Conditional on $(S(t),\boldsymbol{X})$,the conditional variance of $\mu_{yct}$ is also the conditional second moment, meaning that its variance can be used to describe the magnitude of $\mu_{yct}(\boldsymbol{X},\boldsymbol{U})$. Since $\text{Var}(Y(t)\mid S(t),\boldsymbol{X})=\sigma_{y}^{2}+\text{Var}(\mu_{yct}\mid S(t),\boldsymbol{X})$, and $\text{Var}(Y(t)\mid S(t),\boldsymbol{X})$ is identifiable under standard assumptions, the residual variance $\sigma_{y}^{2}$ can determine the magnitude of $\text{Var}(\mu_{yct}\mid S(t),\boldsymbol{X})$. We can use the residual variance to control the contribution of the unobserved intermediate, with higher $\sigma_{y}^{2}$ indicating less variability of $\mu_{yct}$ and lower $\sigma_{y}^{2}$ indicating greater variability of $\mu_{yct}$. It is not clear what a reasonable bound for $\sigma_{y}^{2}$ is, however, we can use an extension of Assumption 7 to the nonparametric setting to guide this choice. Specifically, we can let $\eta_{A\sim B\mid C}^{2}$ represent nonparametric partial $R^{2}$ values, which are defined as
$\eta_{A\sim B\mid C}^{2}=\dfrac{\text{Var}(E[A\mid B,C])-\text{Var}(E[A\mid C])}{\text{Var}(A)-\text{Var}(E[A\mid C])}.$
We then make the following assumption on the contribution of the unobserved intermediate.
Assumption 8 (Generalized Dominant Observed Effect).
Assume
$\eta^{2}_{Y\sim S_{mis}\mid S_{obs},T,\boldsymbol{X}}\leq\eta^{2}_{Y\sim S_{obs}\mid T,\boldsymbol{X}}.$
Similar to before, this assumption places a restriction on the impact of the unobserved intermediate, which is analogous to restricting the magnitude of $\mu_{yct}$. This assumption implies the following inequality
$\dfrac{(\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}]))^{2}}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}\leq\sigma_{y}^{2}.$ (9)
This shows that we can control the variability of $\mu_{yct}$ through a lower bound on $\sigma_{y}^{2}$, similarly to what we did in the case of parametric, linear models. This is crucially important, because implementing a constraint on the variability of $\mu_{yct}$ is difficult to do in practice with complex models, but a constraint on $\sigma_{y}^{2}$ can easily be implemented within the Bayesian framework through a truncated prior distribution. Unlike in the linear, parametric outcome model, it is not straightforward to derive the implications of this assumption on the resulting size of the partial identification region for causal parameters of interest, though it is expected that the reduction will be large in many instances.
8 Discussion
This paper formalizes inference on principal causal effects when principal ignorability is violated by deriving partial identification results for outcome models when this assumption does not hold, and investigating the implications of this assumption on identification of the crucial association parameter $\rho$. We focused our results on settings with continuous intermediates as this is the scenario that relies most heavily on principal ignorability, though our results also hold for binary intermediates under the same set of assumptions. Our results show that there is an inherent trade-off between the strength of assumptions made and the size of partial identification regions. If principal ignorability, or other assumptions are not made, then all parameters are at best partially identified, and the widths of the partial identification regions for parameters of interest can be exceedingly wide. On the other hand, if strong assumptions are made, then all unknown parameters, including the typically unidentified $\rho$ parameter, can be consistently estimated. We have proposed alternative assumptions that should hold in most applications, which can be applied to sharpen inference and reduce the widths of partial identification regions under mild assumptions. Lastly, many of our results held for linear, parametric outcome models, but we showed that similar ideas can be applied for fully nonparametric outcome models, which are popular in causal inference.
There are a number of areas for future research that could build on this work. One area of interest may be the implementation of Model (8), along with an empirical examination of the extended sign and magnitude issues introduced by violations of principal ignorability. It would also be of interest to study the identification of $\rho$ in this scenario and whether this is possible under certain constraints on the outcome model functions. Additionally, it would be of interest to see if other, plausible assumptions could be used in conjunction with those seen in this manuscript to sharpen inference on partial identification regions. For example, we have not made the monotonicity assumption throughout because it is not sufficient for identification with continuous intermediates, though it is plausible in many applications. It may provide significant information on the distribution of $\boldsymbol{U}$ that could lead to smaller partial identification regions, or more efficient estimation of $\rho$ when it is identified.
Appendix A Proofs
A.1 Proofs Relating to Model (2)
Consider the following reparameterization for the outcome model in (1):
$\begin{split}E[Y_{i}(t)\mid\boldsymbol{U}_{i},\boldsymbol{X}_{i},T_{i}=t]&=\boldsymbol{\beta}_{t}^{T}(\boldsymbol{U}_{i}-(\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i},\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i})^{T})+\lambda_{t}+\boldsymbol{\gamma}^{T}_{t}\boldsymbol{X}_{i}.\end{split}$
where $\boldsymbol{\gamma}_{t}=\boldsymbol{\gamma}+(\beta_{t0}+\beta_{t1})\boldsymbol{\alpha}$. Given $T_{i}=t$, consider following change of variable:
$\begin{split}Y_{i}^{\prime}(t)&=Y_{i}(t)-\boldsymbol{\gamma}_{t}^{T}\boldsymbol{X}_{i},\\ S^{\prime}_{i}(t)&=S_{i}(t)-\boldsymbol{\alpha}^{T}\boldsymbol{X}_{i},\qquad t=0,1.\end{split}$ (10)
An important benefit of (10) is that $(\boldsymbol{\gamma}_{0},\boldsymbol{\gamma}_{1},\boldsymbol{\alpha})$ are identifiable in model (2). A short explanation is as follows: Apparently, $\boldsymbol{\alpha}$ is identifiable. Since
$\begin{split}E[Y_{i}(t)|S^{\prime}_{i}(t),\boldsymbol{X}_{i},T_{i}=t]&=\lambda_{t}+\beta_{t,1-t}E[S^{\prime}_{i}(1-t)|S^{\prime}(t),T_{i}=t]+\beta_{tt}S^{\prime}(t)+\boldsymbol{\gamma}_{t}^{T}\boldsymbol{X}_{i},\end{split}$
$\boldsymbol{\gamma}_{t}$ is also identifiable for $t=0,1$.
Then,
$\begin{split}E[Y^{\prime}_{i}\mid\boldsymbol{U}^{\prime}_{i},\boldsymbol{X}_{i},T_{i}=t]&=\boldsymbol{\beta}_{t}^{T}\boldsymbol{U}^{\prime}_{i}+\lambda_{t},\\ E[\boldsymbol{U}^{\prime}_{i}\mid\boldsymbol{X}_{i}]&=(\phi_{0},\phi_{1})^{T}.\end{split}$
For proofs of Theorems 1 and 2, since $(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{1},\boldsymbol{\alpha},\boldsymbol{\gamma})$ all are assumed known, then we can apply (10). As for proofs of Propositions 1, 2 and 3, since $(\boldsymbol{\gamma}_{0},\boldsymbol{\gamma}_{1},\boldsymbol{\alpha})$ are identifiable, then we can also apply (10).
A.1.1 Proof of Theorem 1
Proof.
We apply (10) and let $Y^{\prime}_{i}=Y^{\prime}_{i}(T_{i})$ and $S^{\prime}_{i}=S^{\prime}_{i}(T_{i})$. Throughout the proof, we always condition on $\boldsymbol{X}$ and $T$, unless specified.
The conditional marginal pdf of $(Y^{\prime}_{i},S^{\prime}_{i})$ is
$(Y^{\prime}_{i},S^{\prime}_{i})\mid\boldsymbol{X}_{i},T_{i}=t\sim N\left((\mu^{\prime}_{yt},\phi_{t}),\begin{pmatrix}\zeta_{t}&\psi_{t}\sigma_{st}\\ \psi_{t}\sigma_{st}&\sigma^{2}_{st},\end{pmatrix}\right),$ (11)
where $\mu^{\prime}_{yt}=\boldsymbol{\beta}_{t}^{T}\boldsymbol{\phi}_{i}+\lambda_{t}$, $\zeta_{t}=\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t}$, and $\psi_{t}=\sigma_{st}\beta_{tt}+\rho\sigma_{s,1-t}\beta_{t,1-t}$.
The conditional marginal pdf of $(Y^{\prime}_{1},\dots,Y^{\prime}_{n},S^{\prime}_{1},\dots,S^{\prime}_{n})$ given $\boldsymbol{X}_{i}=\boldsymbol{x}_{i}$ and $T_{i}=t_{i}$ is
$p(\cdot)=\prod_{i=1}^{n}\bigg{(}(2\pi\sigma_{s,t_{i}})^{-1}(\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg{(}-\frac{1}{2}\bigg{(}\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+\frac{s_{i}^{\prime\prime 2}}{\sigma_{s,t_{i}}^{2}}\bigg{)}\bigg{)}\bigg{)},$
where $y^{\prime\prime}_{i}=y^{\prime}_{i}-\lambda_{t_{i}}-\boldsymbol{\mu}_{s}^{T}\boldsymbol{\beta}_{t_{i}}$, $s^{\prime\prime}_{i}=s^{\prime}_{i}-\phi_{t_{i}}$, and $\zeta_{t_{i}}-\psi_{t_{i}}^{2}=\sigma_{y}^{2}+(1-\rho^{2})\sigma_{s,1-t_{i}}^{2}\beta_{t_{i},1-t_{i}}^{2}$.
Then, the unnormalized log posterior for $\rho$ is
$q_{n}(\rho)=-\frac{1}{2}\sum\limits_{i=1}^{n}\left(\log(\zeta_{t_{i}}-\psi_{t_{i}}^{2})+\frac{(y_{i}^{\prime\prime}-\psi_{t_{i}}s_{i}^{\prime\prime}/\sigma_{s,t_{i}})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}+s_{i}^{\prime\prime 2}/\sigma_{s,t_{i}}^{2}\right)+\log(f(\rho)),$
where $f(\rho)$ is the prior for $\rho$.
Let
$l_{i}(\rho\mid t)=-\frac{1}{2}\left(\log(\zeta_{t}-\psi_{t}^{2})+\frac{(y_{i}^{\prime\prime}-\psi_{t}s_{i}^{\prime\prime}/\sigma_{st})^{2}}{(\zeta_{t}-\psi_{t}^{2})}+s_{i}^{\prime\prime 2}/\sigma_{st}^{2}\right),\qquad t=0,1.$
Then, $q_{n}(\rho)=\sum\limits_{i=1}^{n}(t_{i}l_{i}(\rho\mid 1)+(1-t_{i})l_{i}(\rho\mid 0))+\log(f(\rho))$. Let $n_{0}=\sum\limits_{i=1}^{n}(1-t_{i})$ and $n_{1}=\sum\limits_{i=1}^{n}t_{i}$. By the positivity assumption, $n_{0}\to\infty$ and $n_{1}\to\infty$, as $n\to\infty$. Within the control group ($T_{i}=0$) or the treatment group ($T_{i}=1$), $(Y^{\prime}_{i},S^{\prime}_{i})$ are i.i.d.
Let $\hat{\rho}_{0}=\operatorname*{arg\,max}_{\rho}(\sum\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0})$ and $\hat{\rho}_{1}=\operatorname*{arg\,max}_{\rho}(\sum\limits_{i=1}^{n}t_{i}l_{i}(\rho\mid 1)/n_{1})$. Next, we show that $\hat{\rho}_{0}\to{{\rho}^{\ast}}$ and $\hat{\rho}_{1}\to{{\rho}^{\ast}}$ in probability as $n\to\infty$. By Proposition 7.1 in [14], we need to show that they satisfy the identification condition and the uniform convergence condition within both groups.
For the control group,
  • Identification condition:
    The conditional mean of $l_{i}(\rho\mid 0)$ is
    $\begin{split}E(l_{i}(\rho\mid 0))&=-\frac{1}{2}\left(\log(\zeta_{0}-\psi_{0}^{2})+\frac{(1-p)(\psi_{0}^{2}-2\psi_{0}{{\psi}^{\ast}}_{0}+{{\zeta}^{\ast}}_{0})}{\zeta_{0}-\psi_{0}^{2}}\right),\end{split}$
    Taking derivative w.r.t $\rho$, we have
    $\begin{split}\frac{\partial E(l_{i}(\rho\mid 0))}{\partial\rho}&=\frac{2\sigma_{s1}^{2}\beta_{01}^{2}}{\zeta_{0}-\psi_{0}^{2}}\left(\frac{-\sigma_{s1}^{2}\beta_{01}^{2}\rho^{2}}{\zeta_{0}-\psi_{0}^{2}}-1\right)(\rho-\rho^{*}).\end{split}$
    As long as $\beta_{01}\neq 0$, we have $02\sigma_{s1}^{2}\beta_{01}^{2}/(\zeta_{0}-\psi_{0}^{2})(-\sigma_{s1}^{2}\beta_{01}^{2}\rho^{2}/(\zeta_{0}-\psi_{0}^{2})-1)<$ 0, where $\zeta_{0}-\psi_{0}^{2}=\sigma_{y}^{2}+(1-\rho^{2})\beta_{01}^{2}\sigma_{s1}^{2}>0$. Thus, $E(l_{i}(\rho\mid 0))$ is strictly increasing on $[-1,{{\rho}^{\ast}})$ and strictly decreasing on $({{\rho}^{\ast}},1]$, indicating that $E(l_{i}(\rho\mid 0))$ is uniquely maximized on $[-1,1]$ at ${{\rho}^{\ast}}$.
  • Uniform convergence:
    Note that $\text{Var}\left(\sum\limits_{i=1}^{n}(1-t_{i})l_{i}(\rho\mid 0)/n_{0}\right)=\text{Var}(l_{i}(\rho\mid 0))/n_{0}=4/n_{0}$. By Chebyshev’s inequality, uniform convergence holds.
For the treatment group, similarly, we can show that these also hold.
Since $q_{n}({\rho})=\sum\limits_{i=1}^{n}\left(t_{i}l_{i}(\rho\mid 1)/n_{0}(n_{0}/n)+(1-t_{i})l_{i}(\rho\mid 0)/n_{1}(n_{1}/n)\right)+\log(f(\rho))/n$, the above provides intuition that $\hat{\rho}_{B}$ (posterior mode) converges to ${{\rho}^{\ast}}$.
To complete our proof rigorously, we will show that for any $\epsilon>0$, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$ as $n\to\infty$. If $f(\rho)$ is continuous or bounded, then $\log(f(\rho))/n\to 0$ uniformly as $n\to\infty$. Since $\sum\limits_{t_{i}=t}l_{i}(\rho\mid t)/n_{t}$ uniformly converges to $E(l_{i}(\rho\mid t))$, we have $|q_{n}({\rho})-E(l_{i}(\rho\mid 0))(n_{0}/n)-E(l_{i}(\rho\mid 1))(n_{1}/n)|\to 0$ uniformly in probability, and $P(q_{n}(\rho)\geq q_{n}({{\rho}^{\ast}}):|\rho-{{\rho}^{\ast}}|>\epsilon)\to 0$, that is, $P(|\hat{\rho}_{B}-{{\rho}^{\ast}}|>\epsilon)\to 0$.
A.1.2 Proof of Theorem 2
Proof.
We inherit the notations from the proof of Theorem 1. Throughout the proof, we always condition on $\boldsymbol{X}$ and $T$, unless specified. In here we using the Laplace approximation to obtain the asymptotic approximation of posterior variance. Note that the Bernstein–von Mises Theorem provide the justifications for the Laplace approximation.
$\begin{split}E(\rho\mid\mathcal{D})&\approx\frac{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B},\end{split}$
where $q_{n}^{\prime\prime}=\dfrac{\partial^{2}q_{n}(\rho)}{\partial\rho^{2}}$.
Similarly,
$\begin{split}E(\rho^{2}\mid\mathcal{D})&\approx\frac{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\rho^{2}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}{\int_{\rho^{*}-\delta_{1}}^{\rho^{*}+\delta_{2}}\exp(q_{n}(\hat{\rho}_{B})+\frac{1}{2}q_{n}^{\prime\prime}(\hat{\rho}_{B})(\rho-\hat{\rho}_{B})^{2}+o((\rho-\hat{\rho}_{B})^{2})d\rho}\\ &=\hat{\rho}_{B}^{2}-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}.\end{split}$
Therefore, an approximation of posterior variance is $-\frac{1}{q_{n}^{\prime\prime}(\hat{\rho}_{B})}$.
By the Strong Law of Large Number (SLLN), we have
$\sum\limits_{t_{i}=t}l^{\prime\prime}_{i}(\rho\mid t)n_{t}^{-1}\to E(l^{\prime\prime}_{i}(\rho\mid t)),\text{ almost surely,}$,
where $l^{\prime\prime}_{i}(\rho\mid t)=\frac{\partial^{2}l_{i}(\rho|t)}{\partial\rho^{2}}$.
Since $q^{\prime\prime}_{n}(\rho)/n=\sum\limits_{t_{i}=0}(l^{\prime\prime}_{i}(\rho\mid 0)/n_{0})(n_{0}/n)+\sum\limits_{t_{i}=1}(l^{\prime\prime}_{i}(\rho\mid 1)/n_{1})(n_{1}/n)$, then an asymptomatic approximation of $q_{n}^{\prime\prime}(\rho)/n$ can be
$\begin{split}q^{\prime\prime}_{n}(\rho)/n\approx E(l^{\prime\prime}_{i}(\rho\mid 0))(1-\bar{T})+E(l^{\prime\prime}_{i}(\rho\mid 1))\bar{T},\end{split}$
where $\bar{T}=n_{1}/n$.
Note that there is only a constant difference (w.r.t $\rho$) between $l^{\prime\prime}_{i}(\rho\mid t)$ and the log-likelihood of $(Y^{\prime}_{t},S^{\prime}_{t})$. Applying the second Bartlett’s Identity to it, we have
$\begin{split}Var(\rho\mid\mathcal{D})&\sim n^{-1}\left(\bar{T}E\bigg{(}(\frac{\partial l_{i}(\rho|1)}{\partial\rho})^{2}\bigg{)}+(1-\bar{T})E\bigg{(}(\frac{\partial l_{i}(\rho|0)}{\partial\rho})^{2}\bigg{)}\right)^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}\\ &\sim n^{-1}\bigg{(}\bar{T}\beta_{10}^{2}\sigma_{s0}^{2}\bigg{(}\frac{2\rho^{2}\beta_{10}^{2}\sigma_{s0}^{2}}{(\zeta_{1}-\psi_{1}^{2})^{2}}+\frac{1}{\zeta_{1}-\psi_{1}^{2}}\bigg{)}+\\ &\qquad(1-\bar{T})\beta_{01}^{2}\sigma_{s1}^{2}\bigg{(}\frac{2\rho^{2}\beta_{01}^{2}\sigma_{s1}^{2}}{(\zeta_{0}-\psi_{0}^{2})^{2}}+\frac{1}{\zeta_{0}-\psi_{0}^{2}}\bigg{)}\bigg{)}^{-1}\bigg{|}_{\rho=\hat{\rho}_{B}}.\end{split}$ (12)
Since the right-hand side of (12) is a continuous function of $\rho$, and $\hat{\rho}_{B}$ is consistent for ${{\rho}^{\ast}}$, by the continuous mapping theorem, we can replace $\hat{\rho}_{B}$ with ${{\rho}^{\ast}}$ asymptotically. ∎
For Proposition 1, 2 and 3, we present the proof considering covariates.
A.1.3 Proof of Proposition 1
Proof.
We apply (10) and let $Y^{\prime}_{i}=Y^{\prime}_{i}(T_{i})$ and $S^{\prime}_{i}=S^{\prime}_{i}(T_{i})$. For the marginal parameters in (11), we have
$\begin{cases}&\mu_{y0}^{\prime}=\lambda_{0}+\beta_{00}\phi_{0}+\beta_{01}\phi_{1}\\ &\mu_{y1}^{\prime}=\lambda_{1}+\beta_{10}\phi_{0}+\beta_{11}\phi_{1}\\ &\zeta_{0}=\sigma_{y}^{2}+\boldsymbol{\beta}_{0}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{0}\\ &\zeta_{1}=\sigma_{y}^{2}+\boldsymbol{\beta}_{1}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{1}\\ &\psi_{0}=\sigma_{s0}\beta_{00}+\rho\sigma_{s1}\beta_{01}\\ &\psi_{1}=\sigma_{s1}\beta_{11}+\rho\sigma_{s0}\beta_{10}\end{cases}$ (13)
The general solutions of (13) is
$\begin{cases}&\beta_{t,1-t}^{2}=(\zeta_{t}-\psi_{t}^{2}-\sigma_{y}^{2})/(1-\rho^{2})/\sigma^{2}_{s,1-t}\\ &\beta_{tt}=(\psi_{t}-\rho\sigma_{s,1-t}\beta_{t,1-t})/\sigma_{st}\\ &\lambda_{t}=\mu^{\prime}_{yt}-(\beta_{t0}\phi_{0}+\beta_{t1}\phi_{1}).\end{cases}$ (14)
In (14), except for signs, there is only one unknown parameter $\sigma_{y}^{2}$.
Since $\beta_{t,1-t}^{2}\geq 0$, (14) also implies that $0\leq\sigma_{y}^{2}\leq\min\limits_{t=0,1}(\zeta_{t}-\psi_{t}^{2})$. Plugging this in first equation of (14), we have
$\frac{\zeta_{t}-\psi_{t}^{2}-\min\limits_{t=0,1}(\zeta_{t}-\psi_{t}^{2})}{(1-\rho^{2})\sigma^{2}_{s,1-t}}\leq\beta_{t,1-t}^{2}\leq\frac{\zeta_{t}-\psi_{t}^{2}}{(1-\rho^{2})\sigma^{2}_{s,1-t}}.$
Since $\sigma_{s1}^{2}\beta_{01}^{2}\leq\sigma_{s0}^{2}\beta_{10}^{2}$ (it is equivalent to $\text{Var}(Y(1)\mid S(1),\boldsymbol{X})\geq\text{Var}(Y(0)\mid S(0),\boldsymbol{X})$), then it can be simplified as
$\begin{split}0\leq&\beta_{01}^{2}\leq\frac{\zeta_{0}-\psi_{0}^{2}}{(1-\rho^{2})\sigma^{2}_{s1}},\\ \frac{\zeta_{1}-\psi_{1}^{2}-(\zeta_{0}-\psi_{0}^{2})}{(1-\rho^{2})\sigma^{2}_{s0}}\leq&\beta_{10}^{2}\leq\frac{\zeta_{1}-\psi_{1}^{2}}{(1-\rho^{2})\sigma^{2}_{s0}}.\end{split}$
Note that (14) also requires that
$\sigma_{s0}^{2}\beta_{10}^{2}-\sigma_{s1}^{2}\beta_{01}^{2}=\frac{\zeta_{1}-\psi_{1}^{2}-(\zeta_{0}-\psi_{0}^{2})}{1-\rho^{2}}.$
We plug in $\text{Var}(Y(t)\mid S(t),\boldsymbol{X})=\text{Var}(Y(t)\mid S(t),\boldsymbol{X})=\zeta_{t}-\psi_{t}^{2}$ and $\text{Var}(S(t)\mid\boldsymbol{X})=\sigma_{st}^{2}$ to obtain the desired forms. ∎
A.1.4 Derivation for (4)
$\begin{split}PCE(\boldsymbol{u})&=E(PCE(\boldsymbol{u},\boldsymbol{x})\mid\boldsymbol{U}=\boldsymbol{u})\\ &=E((\beta_{10}-\beta_{00})s_{0}+(\beta_{11}-\beta_{01})s_{1}+\lambda_{1}-\lambda_{0}\mid\boldsymbol{U}=\boldsymbol{u})\\ &=(\beta_{10}-\beta_{00})s_{0}+(\beta_{11}-\beta_{01})s_{1}+\lambda_{1}-\lambda_{0}\end{split}$
Using (14), we first replace $\lambda_{t}$ and then replace $\beta_{tt}$. After simplification, we have
$\begin{split}PCE(\boldsymbol{u})&=(\beta_{10}-\frac{\psi_{0}}{\sigma_{s0}}+\frac{\sigma_{s1}}{\sigma_{s0}}\rho\beta_{01})(s_{0}-\phi_{0})+\\ &\qquad(\frac{\psi_{1}}{\sigma_{s1}}-\frac{\sigma_{s0}}{\sigma_{s1}}\rho\beta_{10}-\beta_{01})(s_{1}-\phi_{1})+(\mu^{\prime}_{y1}-\mu^{\prime}_{y0}),\end{split}$
A.1.5 Proof of Proposition 2
Proof.
By directly applying Assumption 6 to Proposition 1, we can obtain Proposition 2. ∎
A.1.6 Proof of Proposition 3
Proof.
Assumption 7 is equivalent to
$\begin{split}\frac{(\text{Cov}(Y(t),S(1-t)\mid S(t),\boldsymbol{X}))^{2}}{\text{Var}(S(1-t)\mid S(t),\boldsymbol{X})\text{Var}(Y\mid S(t),\boldsymbol{X})}&\leq\frac{(\text{Cov}(Y(t),S(t)\mid\boldsymbol{X}))^{2}}{\text{Var}(Y\mid\boldsymbol{X})\text{Var}(S(t)\mid\boldsymbol{X})}\\ \frac{\beta_{t,1-t}^{2}\sigma^{4}_{s,1-t}(1-\rho^{2})^{2}}{\sigma^{2}_{s,1-t}(1-\rho^{2})(\sigma_{y}^{2}+\beta_{t,1-t}^{2}\sigma^{2}_{s,1-t}(1-\rho^{2}))}&\leq\frac{(\beta_{tt}\sigma_{st}+\rho\beta_{t,1-t}\sigma_{s,1-t})^{2}}{\sigma_{st}^{2}(\sigma_{y}^{2}+\boldsymbol{\beta}_{t}^{T}\boldsymbol{\Sigma}_{s}\boldsymbol{\beta}_{t})}\\ \end{split}$
Using (14), we have
$\begin{split}\frac{(\zeta_{t}-\psi_{t}^{2})^{2}}{\zeta_{t}}&\leq\sigma_{y}^{2}.\end{split}$
Combing it with the proof for Proposition 1, the corresponding partial identification region can be obtained.
Note that the above proof also shows that Assumption 7 is equivalent to
$\dfrac{(\text{Var}(Y(t)\mid S(t),\boldsymbol{X}))^{2}}{\text{Var}(Y(t)\mid\boldsymbol{X})}\leq\text{Var}(Y(t)\mid S(t),S(1-t),\boldsymbol{X}).$
A.2 Proofs Relating to Model (8)
A.2.1 Proof of Theorem 3
Proof.
Note that $Y^{\prime}(t)$ and $S^{\prime}(1-t)$ are also used in this proof, but they have different definitions than before.. Let $f_{t}(S(t),S(1-t),\boldsymbol{X})$ denote $\mu_{yt}(\boldsymbol{X},S(t))+\mu_{yct}(\boldsymbol{X},S(t),S(1-t))$. Then
$f_{t}\big{(}S(t),M^{-1}(S(1-t)),\boldsymbol{X}\big{)}=\mu_{yt}\big{(}\boldsymbol{X},S(t)\big{)}+\mu_{yct}\big{(}\boldsymbol{X},S(t),M^{-1}(S(1-t))\big{)},$
and
$Y^{\prime}(t)=f_{t}(S(t),M^{-1}(S(1-t)),\boldsymbol{X})+\epsilon.$
Although $M$ could relate to $S(t)$ and $\boldsymbol{X}$, we do not include them in the $M$ for simplicity of notation. Let $S^{\prime}(1-t)$ denote $M(S(1-t))$ and $p_{\epsilon}$ denote the pdf of $\epsilon$. The the conditional distribution of $Y^{\prime}(t)$ given $S(t)$ and $\boldsymbol{X}$ is
[…]
The third equality holds since $S^{\prime}(1-t)\mid S(t),\boldsymbol{X}$ and $S(1-t)\mid S(t),\boldsymbol{X}$ have the same distribution and we also change the notation from $s_{1-t}$ to $s^{\prime}_{1-t}$. The fourth equality holds since we apply the change of variable $S(1-t)=M^{-1}(S^{\prime}_{1-t})$.
The the conditional distribution of $(Y^{\prime}(t),S(t))$ given $\boldsymbol{X}$ is
$\begin{split}p\big{(}Y^{\prime}(t)=y,S(t)=s_{t}\mid\boldsymbol{X}\big{)}&=p\big{(}Y^{\prime}(t)=y\mid\boldsymbol{X},S(t)=s_{t}\big{)}p\big{(}S(t)=s_{t}\mid\boldsymbol{X}\big{)}\\ &=p\big{(}Y(t)=y\mid\boldsymbol{X},S(t)=s_{t}\big{)}p\big{(}S(t)=s_{t}\mid\boldsymbol{X}\big{)}\\ &=p\big{(}Y(t)=y,S(t)=s_{t}\mid\boldsymbol{X}\big{)}.\end{split}$
Therefore, $(Y^{\prime}(t),S(t))\mid\boldsymbol{X}$ and $(Y(t),S(t))\mid\boldsymbol{X}$ follow the same distribution.
A.2.2 Proof of an equivalent expression for Assumption 8
Proof.
Using the definition of Pearson’s correlation ratio, Assumption 8 is equivalent to:
$\begin{split}&\dfrac{\text{Var}(E[Y(t)\mid\boldsymbol{U},\boldsymbol{X}])-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])}\leq\\ &\qquad\dfrac{\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}.\end{split}$
Since $\epsilon\perp\!\!\!\!\perp E[Y(t)\mid\boldsymbol{U},\boldsymbol{X}]$, then
$\text{Var}(Y(t))=\text{Var}(E[Y(t)\mid\boldsymbol{U},\boldsymbol{X}])+\sigma_{y}^{2}.$
Combining them, we have
$\begin{split}&\dfrac{\text{Var}(Y(t))-\sigma_{y}^{2}-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])}\leq\\ &\qquad\dfrac{\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}])-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}.\end{split}$
We simplify the above, then we have
$\dfrac{(\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid S(t),\boldsymbol{X}]))^{2}}{\text{Var}(Y(t))-\text{Var}(E[Y(t)\mid\boldsymbol{X}])}\leq\sigma_{y}^{2}$
Appendix B Details of MCMC sampling
In this section, we detail the general Gibbs updating steps for Model (2). Note that Model (2) includes covariates. Throughout this section, let $\boldsymbol{\theta}_{y}=(\boldsymbol{\beta}_{0},\boldsymbol{\beta}_{1},\lambda_{0},\lambda_{1},\boldsymbol{\gamma})$ and $\boldsymbol{\theta}_{s}=(\boldsymbol{\phi},\boldsymbol{\alpha})$, both of which are column vectors. The full joint pdf of is:
$\begin{split}&\dfrac{1}{(2\pi)^{\frac{3n}{2}}\sigma_{y}^{n}\sigma_{s0}^{n}\sigma_{s1}^{n}(1-\rho^{2})^{\frac{n}{2}}}\exp\bigg{(}-\frac{1}{2}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2}/\sigma_{y}^{2}+\\ &\qquad(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})^{T}\boldsymbol{\Sigma}_{s}^{-1}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})\bigg{)}\bigg{)},\end{split}$ (15)
where
$\boldsymbol{d}_{i}=\begin{pmatrix}\boldsymbol{u}_{i}(1-t_{i})\\ \boldsymbol{u}_{i}t_{i}\\ 1-t_{i}\\ t_{i}\\ \boldsymbol{x}_{i}\end{pmatrix}\ \text{and}\ \boldsymbol{E}_{i}=(\boldsymbol{e}_{i0},\boldsymbol{e}_{i1})=\begin{pmatrix}1&0\\ 0&1\\ \boldsymbol{x}_{i}&\boldsymbol{x}_{i}\\ \end{pmatrix}.$
We consider conjugate priors for all parameters except $\rho$ as follows:
$\begin{split}&\boldsymbol{\beta}_{t}\sim N(\boldsymbol{\mu}_{\boldsymbol{\beta}_{t}},\Sigma_{\boldsymbol{\beta}_{t}}),\ \lambda_{t}\sim N(\mu_{\lambda_{t}},\sigma_{\lambda_{t}}^{2}),\ \phi_{t}\sim N(\mu_{\phi_{t}},\sigma_{\phi_{t}}^{2}),\ \boldsymbol{\gamma}\sim N(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\Sigma_{\boldsymbol{\gamma}}),\\ &\boldsymbol{\alpha}\sim N(\boldsymbol{\mu}_{\boldsymbol{\alpha}},\Sigma_{\boldsymbol{\alpha}}),\ \sigma_{y}^{2}\sim IG(\eta_{y},\nu_{y}),\ \sigma_{s0}^{2}\sim IG(\eta_{s0},\nu_{s0}),\ \sigma_{s1}^{2}\sim IG(\eta_{s1},\nu_{s1}),\end{split}$
For $\rho$, it has a flat prior.
Then, the updating steps for Gibbs sampling are as follows
  1. a.
    We can update the unobserved values of the intermediate $S(1-t_{i})$ from
    $S(1-t_{i})\mid\cdot\sim N(\mu_{si,mis},\sigma_{si,mis}^{2}),$
    where
    $\sigma_{si,mis}^{2}=\left(\frac{\beta_{t_{i},1-t_{i}}^{2}}{\sigma_{y}^{2}}+\frac{1}{\sigma_{s,1-t_{i}}^{2}(1-\rho^{2})}\right)^{-1},$
    and
    $\begin{split}\mu_{si,mis}&=\left(\frac{\beta_{t_{i},1-t_{i}}^{2}}{\sigma_{y}^{2}}+\frac{1}{\sigma_{s,1-t_{i}}^{2}(1-\rho^{2})}\right)^{-1}\bigg{(}-(\beta_{t_{i},t_{i}}s_{i}+\lambda_{t_{i}}+\boldsymbol{\gamma}^{T}\boldsymbol{x}_{i}-y_{i})\beta_{t_{i},1-t_{i}}/\sigma_{y}^{2}+\\ &\qquad(\phi_{1-t_{i}}+\boldsymbol{\alpha}^{T}\boldsymbol{x}_{i})/\sigma_{s,1-t_{i}}^{2}/(1-\rho^{2})+\rho(s_{i}-\phi_{t_{i}}-\boldsymbol{\alpha}^{T}\boldsymbol{x}_{i})/\big{(}\sigma_{s0}\sigma_{s1}(1-\rho^{2})\big{)}\bigg{)}.\end{split}$
  2. b.
    We can update $\boldsymbol{\theta}_{y}$ from the following conditional distribution
    $\boldsymbol{\theta}_{y}\mid\cdot\sim N\left(\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\bigg{(}\sum\limits_{i=1}^{n}y_{i}\boldsymbol{d}_{i}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{y}}\bigg{)},\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\right),$
    where
    $\boldsymbol{\Sigma}_{\boldsymbol{\theta}_{y}}=\text{diag}(\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{0}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{1}},\sigma^{2}_{\lambda_{0}},\sigma^{2}_{\lambda_{1}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}).$
  3. c.
    We update $\boldsymbol{\theta}_{s}$ from
    $\boldsymbol{\theta}_{s}\mid\cdot\sim N\left(\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{E}_{i}^{T}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\bigg{)}^{-1}\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{u}_{i}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{s}}\bigg{)},\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{E}_{i}^{T}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\bigg{)}^{-1}\right),$
    where $\boldsymbol{\Sigma}_{\boldsymbol{\theta}_{s}}=\text{diag}(\sigma_{\phi_{0}}^{2},\sigma_{\phi_{1}}^{2},\boldsymbol{\Sigma}_{\boldsymbol{\alpha}})$.
  4. d.
    We can update $\sigma_{y}$ from
    $\sigma_{y}^{2}\mid\cdot\sim IG\left(\frac{n}{2}+\eta_{y},\sum\limits_{i=1}^{n}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2})/2+\nu_{y}\right).$
  5. e.
    If we consider $\sigma_{s0}\neq\sigma_{s1}$, then we need to apply Metropolis-Hasting algorithm to update $\sigma_{s0}^{2}$ and $\sigma_{s1}^{2}$ from
    1. (a)
      $\begin{split}p(\sigma_{s0}^{2}\mid\cdot)&\propto\sigma_{s0}^{-n-2-2\eta_{s0}}\exp\bigg{(}-\frac{1}{2(1-\rho^{2})\sigma_{s1}^{2}}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})^{2}\sigma_{s1}^{2}/\sigma_{s0}^{2}-\\ &\qquad 2\rho\sigma_{s1}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s0}\bigg{)}-\frac{\nu_{s0}}{\sigma_{s0}^{2}}\bigg{)}.\end{split}$
    2. (b)
      $\begin{split}p(\sigma_{s1}^{2}\mid\cdot)&\propto\sigma_{s1}^{-n-2-2\eta_{s1}}\exp\bigg{(}-\frac{1}{2(1-\rho^{2})\sigma_{s0}^{2}}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})^{2}\sigma_{s0}^{2}/\sigma_{s1}^{2}-\\ &\qquad 2\rho\sigma_{s0}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s1}\bigg{)}-{\nu_{s1}}/{\sigma_{s1}^{2}}\bigg{)}.\end{split}$
    If we assume $\sigma_{s0}=\sigma_{s1}=\sigma_{s}$, then given a conjugate inverse gamma prior, $\sigma_{s}^{2}\sim IG(\eta_{s},\nu_{s})$, we can update $\sigma_{s}^{2}$ using the following conditional distribution:
    $\begin{split}\sigma_{s}^{2}\mid IG\bigg{(}n+\eta_{s},\sum\limits_{i=1}^{n}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})^{T}\boldsymbol{R}_{s}^{-1}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})/2+\nu_{s}\bigg{)},\end{split}$
    where
[…]
  6. f.
    Instead of updating $\rho$ from posterior of full data including unobserved intermediates, we update $\rho$ using marginal posterior with observed intermediate variables only. Then, we have
    $\begin{split}p(\rho\mid\cdot)&\propto\prod_{i=1}^{n}\bigg{(}(\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg{(}-\frac{1}{2}\sum\limits_{i=1}^{n}\frac{(\psi_{t_{i}}s^{\prime\prime}_{i}/\sigma_{st}-y^{\prime\prime}_{i})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}\bigg{)}\bigg{)},\end{split}$
    where given $T_{i}=t$, $y^{\prime\prime}_{i}=y_{i}-\lambda_{t}-\boldsymbol{\gamma}^{T}\boldsymbol{x}_{i}-\boldsymbol{\mu}_{si}^{T}\boldsymbol{\beta}_{t}$ and $s^{\prime\prime}_{i}=s_{i}-\phi_{t}-\boldsymbol{\alpha}^{T}\boldsymbol{x}_{i}$. Since the conditional distribution of $\rho$ is intractable, we apply a Metropolis-Hastings algorithm to update $\rho$.
If there are no covariates, one can simply remove $\boldsymbol{x}_{i}$ and corresponding parameters $(\boldsymbol{\alpha},\boldsymbol{\gamma})$. For truncated priors, one can modify the above Gibbs updating steps to obtain the truncated versions.
Appendix C Additional plots
We ran additional MCMC chains under the setting described in Section 6.2 with $n$ varying from 300 to 9600. We plotted $\log(n)$ versus the logarithm of the estimated posterior variance, as shown in Fig. 2. This plot shows that when the sample size is large (over 1000), it is nearly linear, indicating that the posterior variance of $\rho$ is approximately $O(n^{-1})$, consistent with the rate stated in Theorem 2.
Figure 2: $\log(n)$ vs. logarithm of the estimated posterior variance